rxcklesskaisher:

Select three ratios that are equivalent to 7:67:67, colon, 6. Choose 3 answers: Choose 3 answers: (Choice A) 12:14 (Choice B, ) 21:18 (Choice C) 42:36 (Choice D) D 63:54 (Choice E) E 84:62

AZ:

I'm guessing the ratio is just 7:6 is that correct?

rxcklesskaisher:

yes

rxcklesskaisher:

i half to pick three of the choses thats equivalent to it

AZ:

okay so look at the answer choices we have a ratio of 7 to 6 for it to be equivalent, try to see which ones would be simplified down to 7:6 let's look at the first one 12:14 if you divide both numbers by 2, what do you get?

rxcklesskaisher:

6 and 2

AZ:

mhmm 6 is correct but what is 14 / 2

rxcklesskaisher:

7*

AZ:

there you go so remember the ratio we're looking for is 7:6 but we just said 12:14 is the same as 6:7 it's not the same order so A isn't correct

AZ:

what about B 21:18 what do you get if you divide both numbers by 3?

rxcklesskaisher:

7 and 6

AZ:

so is that the same ratio as 7 to 6?

rxcklesskaisher:

no

AZ:

why not you just said 21:18 is the same thing as 7:6 when you divide both numbers by 3 so isn't that the same ratio as 7:6 in the original question which we're looking for?

rxcklesskaisher:

it is because 7x7 is 21

AZ:

7*3 = 21 and 6 * 3 = 21 and so if we have 7:6 it's the same ratio as 21:18 because if you multiplied 7 and 6 by 3, you get 21 and 18 does that make sense?

rxcklesskaisher:

yes

AZ:

good, so B is correct what about C. 42:36 what do you get you divide both numbers by 6?

rxcklesskaisher:

yes

AZ:

yes what

rxcklesskaisher:

c

AZ:

Okay, C is also correct what about D? 63:54 divide both numbers by 9 does it also have a ratio of 7:6 ?

rxcklesskaisher:

no

AZ:

are you sure? what is 63/9 = what is 54/9 = is that the same ratio as 7:6

rxcklesskaisher:

no D

AZ:

.. as in D is correct or wrong?

rxcklesskaisher:

correct

AZ:

and that is correct

AZ:

so now you have selected your three correct answers and you don't even need to bother checking that last answer choice